A point charge q1 = + 7.00 x 10-10 C is at the point x = 0.900m, y= 0.700 m and a second point charge q2 = -3.00 x 10-10 C is at the point x = 0.900 m, y= 0m . Calculate the magnitude and direction of the resultant electric field at the origin due to these charges. 



1
Expert's answer
2022-03-29T12:29:47-0400


Electric field due to charge"q_1"


"E_1=\frac{kq_1}{r^2}"

"E_1=\frac{9\times10^9\times7\times10^{-10}}{\sqrt{(1.3})^2}=4.846N/C"

"E_2=\frac{kq_2}{r^2}"


"E_2=-\frac{9\times10^9\times3\times10^{-10}}{{(0.9})^2}=-3.33N/C"

"E_{net}=\sqrt{E_1^2+E_2^2+2E_1E_2cos\phi}"

"\phi=45°"


"E_{net}=\sqrt{4.846^2+(-3.33)^2+2\times4.846\times3.33\times cos135}"

"E_{net}=3.42N\/C"

Direction OP direction according to daigram